OpenStudy (anonymous):

Solve by completing the square. x² + 10x = -9

OpenStudy (anonymous):

x^2+10x+(10/2)^2= -9 + 25 x^2 + 10x + 25 = 16 x^2+10x +9 = 0 (x+9)(x+1)=0

OpenStudy (anonymous):

how did you get 16 ?

OpenStudy (anonymous):

and then to 0 ?

OpenStudy (anonymous):

he added 25 from -9

OpenStudy (anonymous):

its a rule when you are solving complete square

OpenStudy (anonymous):

Oohh I see now and then he subtracted the 25 and the 16

OpenStudy (anonymous):

Would one of them be negative ?

OpenStudy (anonymous):

nooo because once you get the complete polynomial you have to add the 25 to the other side

OpenStudy (anonymous):

But positive 9 and 1 isnt an answer . Its the answer is negative 9 and 1 .

OpenStudy (anonymous):

nooo you have to use the same operator on both sides

OpenStudy (anonymous):

but its wrong on my coursework ? it says it has to be -9 & -1

OpenStudy (anonymous):

x = -9 and -1

OpenStudy (anonymous):

I understand what you guys are saying but Its not wirk in my coursework

OpenStudy (anonymous):

ohhh i see what you are talking about (x+9)(x+1)=0, now you're solving for x so x will be -9 and -1

OpenStudy (anonymous):

Yeah but do you solve it the same way that you did it ?

OpenStudy (anonymous):

yeah that is the process how you would solve for complete square for any problems

OpenStudy (anonymous):

okay cool thank you so much !(: Could you tell me the steps tho not like work them out in a problem but like telling what I do in a simple way :p ? if that makes sense ? sorry haha .

OpenStudy (anonymous):

use the number on the x divide it by 2 and then square it, once you add the number be sure to add to the right side also then just move the numbers on the right and solve for x by factoring

OpenStudy (anonymous):

you got it or you haven't gotten it yet?

OpenStudy (anonymous):

@galacticwavesXX sorta ? i'm working on another problem but I;m stuck . the orignal question is x^2-6x=5

OpenStudy (anonymous):

and so far got x^2+(-6/2)^2=5 , then x^2-6x+9=5?

OpenStudy (anonymous):

yess!! that is correct so far